A firm faces a demand law q = 50-4p where q and p are the quantity and prices of the good respectively. Find out the price at which total revenue is maximized.



1
Expert's answer
2022-01-16T13:26:54-0500

Total revenue is TR = p×q = (12.5 - 0.25q)×q.

Total revenue is maximized, when:

"MR = TR'(q) = 12.5 - 0.5q = 0,"

0.5q = 12.5,

q = 25 units.

p = 12.5 - 0.25×25 = 6.25.